Use the Ksp values to calculate the molar solubility of each of the following compounds in pure water.

MX (Ksp = 5.35
Answer #1

MX <-> M+ and X-

the solubility equilbrium is expressed as:

Ksp = [M+][X-]

5.35*10^-11 = [M+][X-]

[MX] = [M+] = [X-] = s

Since for 1 mol of MX, you dissolve 1 mol of M+ which is equal to 1 mol of X- in solution... Let us call these "solubility"

Substitute:

5.35*10^-11 = s*s

s^2 =5.35*10^-11

s = sqrt(5.35*10^-11 ) = 7.3*10^-6

Solubility of MX = 7.3*10^-6 mol/liter